A call option on Project Cash Flow Consulting Inc.'s stock (PCF) has a market price of $7.00.
The stock currently sells for $30/share (P), and the option has a strike price of $25/share (X).
What is the exercise value (payoff) of the call option?
What is the option's time value based on the option's market price and the exercise value?
SHOW WORK HERE, HIGHLIGHT FINAL ANSWER IN YELLOW
Answer #1

Payoff of Call Option = Max(0,(Spot Price - Strike Price))

Payoff of Call Option = (30 - 25)

Payoff of Call Option = $5

Time Value = Premium of Option - Intrinsic value

Time Value = 7 - 5

Time Value = $2